# Dirichlet-to-Neumann operator for the Helmholtz problem with general wavenumbers on the $n$-sphere

### v1

@@ 3.2 The DtN operator for the Helmholtz problem on a sphere
Hence (3.8) results with $C_{\operatorname*{tr}}^{2}=3(1+\max\{1,\mathopen{}\mathclose{{}\left(1+C_{\omega,1}^{2}}\right)C_{\operatorname*{F}}^{2}\})C_{\Omega,0}^{2}/2$ in
@@ 4 Spectral expansion of the DtN operator
The $\operatorname*{DtN}$ operator has an explicit series expansion in $n$-dimensional polar coordinates in terms of Laplace-Beltrami eigenfunctions known for $s=\operatorname{i}\mathbb{R}$, e.g., from [2] for $n=2$ and [14, Eqn. (2.6.92)] for $n=3$ dimensions. An extension to general wavenumbers $s\in\mathbb{C}_{\geq 0}$ and dimensions $n\geq 2$ can be derived from [12, Chap. 9].
@@ 4.1 Spectral representation of the Helmholtz DtN
The spherical part of the DtN operator is expanded by an orthonormal basis of the eigenfunctions of the Laplace-Beltrami operator, whose properties are briefly recalled, e.g., from [16, Sec. 22.2]. The (negative) Laplace-Beltrami $-\Delta_{S_{1}}$ on the unit sphere $S_{1}\in\mathbb{R}^{n}$ is defined by
@@
and that the eigenfunctions are chosen such that $\{Y_{m,j}\}_{j\in J_{m},m\in\mathbb{N}_{0}}$ forms a orthonormal basis of $L^{2}\mathopen{}\mathclose{{}\left(S_{1}}\right)$. Explicitly it holds [16, Thm. 22.1] that
@@
The ODE (4.4) is closely related to the (modified) Bessel’s equation (see, e.g., [12, p. 278] for details). For any $s\in\mathbb{C}^{\bullet}_{\geq 0}$, a fundamental system is given by $f_{m,\nu}^{(1)}\mathopen{}\mathclose{{}\left(sr}\right)$ and $f_{m,\nu}^{(2)}\mathopen{}\mathclose{{}\left(sr}\right)$ with
@@ 4.2 Analysis of the spectral coefficients
A holomorphic extension of $M_{\mu}^{2}$ in (4.14) to $\mathbb{C}_{>0}$ is known from Nicholson’s integral formula [18, §13.73] as
@@
Step 1 departs with general remarks. The modified Bessel function $K_{m+\nu}=K_{\mu}$ has no zeroes [18, pp. 511–513] in $\mathbb{C}_{\geq 0}$ and is holomorphic [4, §10.25.ii] in $\mathbb{C}^{\bullet}_{\geq 0}$. Hence $z_{m,\nu}\mathopen{}\mathclose{{}\left(s}\right)$ is well defined and holomorphic for $s\neq 0$. The product rule and the recurrence relations [4, §10.29.2] provide
@@
The investigation of $w_{\mu}(k)$ departs with the decomposition into its real and imaginary parts known from [13, p. 1877] and is repeated here for convenience. First observe
@@
by expanding (4.24) with $H_{\mu}^{(2)}(k)=J_{\mu}(k)-\operatorname{i}Y_{\mu}(k)$. The Wronskian identity $J_{\mu}(k)Y_{\mu}^{\prime}(k)-Y_{\mu}(k)J_{\mu}^{\prime}(k)=2/(\pi k)$ from [4, §10.5.2] is a consequence of the recurrence relations for Bessel functions. This reveals the characterisation [13, Eqn. (3.10)] of $w_{\mu}$, namely
@@
Since $kM_{\mu}^{2}(k)$ is non-increasing for $\mu\geq 1/2$ and converges to $2/\pi$ as $k\rightarrow\infty$ [18, §13.74], one has $2\leq\pi kM_{\mu}^{2}(k)<\infty$ and
@@
The situation is different for $\mu=0$ where $M_{0}^{2}(k)$ admits no lower bound for $k\rightarrow 0$ that is linear in $1/k$. However, the asymptotic expansion [18, §13.75] of $M_{0}^{2}(k)$ reveals the convergence $kM_{0}^{2}(k)\rightarrow 2/\pi$ as $k\rightarrow\infty$ with the lower bound (see [4, §10.18.iii])
@@ 4.3 Spectral representation of the Laplace DtN
4.4 Proof of Theorem 3.4
@@
$\mathopen{}\mathclose{{}\left\langle\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)g,\overline{h}}\right\rangle_{S_{1}}=\sum_{m\in\mathbb{N}_{0}}\sum_{j\in J_{m}}z_{m,\nu}(s)\,\widehat{g}_{m,j}\,\overline{\widehat{h}_{m,j}}.$ (4.41)
Since $z_{m,\nu}(s)=\overline{z_{m,\nu}(\overline{s})}$ by Theorem 4.1.i, this proves Theorem 3.4.i. Recall from the definition of the Fourier coefficients (4.11) and the equivalence of norms [10, Rem. 7.6] in $H^{1/2}(S_{1})$ that there exists a constant $C_{1}(n)>0$ that exclusively depends on $n$ such that
@@
The proof of Proposition 3.2 follows mainly from the theory in [12], where integral equation techniques are used. The key role is played by Green’s representation formula and we sketch the arguments in the following. The starting point is the exterior problem [12, Eqn. (7.31)] ( for $\mathcal{P}=-\Delta+s^{2}$, $f=0$, and $\Gamma=S_{R}$ therein) written as:
@@
Here $\mathcal{M}u=0$ is a condition for the behaviour at infinity which is defined in [12, (7.29)]. Theorems 8.9 and 9.6 in [12] translate this condition for our setting to
@@
If the exterior problem (3.1) with condition $\operatorname*{IC}\nolimits_{s}^{\operatorname*{Green}}$ has a solution $u$, then this solution satisfies Green’s representation theorem by [12, Thm. 7.15], namely
@@
where $\operatorname*{SL}\mathopen{}\mathclose{{}\left(s}\right)$ and $\operatorname*{DL}\mathopen{}\mathclose{{}\left(s}\right)$ are the single and double layer potential operators for the operator $\mathopen{}\mathclose{{}\left(-\Delta+s^{2}}\right)$ (see, e.g., [12, Eqn. (7.1)] for details) and $\psi$ is a solution of
@@
with the abbreviations $S:=\frac{1}{2}\mathopen{}\mathclose{{}\left(\gamma^{+}\operatorname*{SL}\mathopen{}\mathclose{{}\left(s}\right)+\gamma^{-}\operatorname*{SL}\mathopen{}\mathclose{{}\left(s}\right)}\right)$ and $T:=\frac{1}{2}\mathopen{}\mathclose{{}\left(\gamma^{+}\operatorname*{DL}\mathopen{}\mathclose{{}\left(s}\right)+\gamma^{-}\operatorname*{DL}\mathopen{}\mathclose{{}\left(s}\right)}\right)$. Note that for any $g\in H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right)$ and $h\in H^{-1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right)$, $\operatorname*{DL}\mathopen{}\mathclose{{}\left(s}\right)g$ and $\operatorname*{SL}\mathopen{}\mathclose{{}\left(s}\right)h$ define functions in $\mathbb{R}^{n}\backslash S_{R}$ and $\gamma^{+}$ denotes the trace from the outer domain $B_{R}^{+}$, while $\gamma^{-}$ is the trace from the inner domain $B_{R}$. It is proved in [12, Theorem 6.11] that the mappings
@@
are continuous linear operators. In this way, well-posedness of (3.1) with condition $\operatorname*{IC}\nolimits_{s}^{\operatorname*{Green}}$ follows if $S$ in (A.1) has a bounded inverse. Theorem 7.6 in [12] implies that $S$ is a Fredholm operator of index 00 so that uniqueness implies bounded invertibility. Uniqueness follows for $n\geq 2$ and $s\in\mathbb{C}^{\bullet}_{\geq 0}$ from [12, Theorem 9.11], for $n\geq 3$ and $s=0$ from [12, Corollary 8.11] and for $n=2$, $s=0$, and $R\neq 1$ from [12, Theorem 8.16.(ii)]. Hence (3.1) with condition $\operatorname*{IC}\nolimits_{s}^{\operatorname*{Green}}$ is well-posed in these three cases (with continuous solution operator). This finishes the proof of Proposition 3.2.1a and 3.2.2b, whereas 3.2.1b directly follows from the implication (4.10).
It remains to prove Proposition 3.2.2a. Uniqueness of solutions is guaranteed by [12, Thm 8.10]. Existence and well-posedness follows from the expansion (4.38) of the solution

### v2

@@ 3.2 The DtN operator for the Helmholtz problem on a sphere
Hence (3.8) results with $C_{\operatorname*{tr}}^{2}=3\mathopen{}\mathclose{{}\left(1+\max\mathopen{}\mathclose{{}\left\{1,\mathopen{}\mathclose{{}\left(1+C_{\omega,1}^{2}}\right)C_{\operatorname*{F}}^{2}}\right\}}\right)C_{\Omega,0}^{2}/2$ in
@@ 4 Spectral expansion of the DtN operator
The $\operatorname*{DtN}$ operator has an explicit series expansion in $n$-dimensional polar coordinates in terms of Laplace-Beltrami eigenfunctions known for $s=\operatorname{i}\mathbb{R}$, e.g., from [2] for $n=2$ and [15, Eqn. (2.6.92)] for $n=3$ dimensions. An extension to general wavenumbers $s\in\mathbb{C}_{\geq 0}$ and dimensions $n\geq 2$ can be derived from [13, Chap. 9].
@@ 4.1 Spectral representation of the Helmholtz DtN
The spherical part of the DtN operator is expanded by an orthonormal basis of the eigenfunctions of the Laplace-Beltrami operator, whose properties are briefly recalled, e.g., from [17, Sec. 22.2]. The (negative) Laplace-Beltrami $-\Delta_{S_{1}}$ on the unit sphere $S_{1}\in\mathbb{R}^{n}$ is defined by
@@
and that the eigenfunctions are chosen such that $\{Y_{m,j}\}_{j\in J_{m},m\in\mathbb{N}_{0}}$ forms a orthonormal basis of $L^{2}\mathopen{}\mathclose{{}\left(S_{1}}\right)$. Explicitly it holds [17, Thm. 22.1] that
@@
The ODE (4.4) is closely related to the (modified) Bessel’s equation (see, e.g., [13, p. 278] for details). For any $s\in\mathbb{C}^{\bullet}_{\geq 0}$, a fundamental system is given by $f_{m,\nu}^{(1)}\mathopen{}\mathclose{{}\left(sr}\right)$ and $f_{m,\nu}^{(2)}\mathopen{}\mathclose{{}\left(sr}\right)$ with
@@ 4.2 Analysis of the spectral coefficients
A holomorphic extension of $M_{\mu}^{2}$ in (4.14) to $\mathbb{C}_{>0}$ is known from Nicholson’s integral formula [19, §13.73] as
@@
Step 1 departs with general remarks. The modified Bessel function $K_{m+\nu}=K_{\mu}$ has no zeroes [19, pp. 511–513] in $\mathbb{C}_{\geq 0}$ and is holomorphic [4, §10.25.ii] in $\mathbb{C}^{\bullet}_{\geq 0}$. Hence $z_{m,\nu}\mathopen{}\mathclose{{}\left(s}\right)$ is well defined and holomorphic for $s\neq 0$. The product rule and the recurrence relations [4, §10.29.2] provide
@@
The investigation of $w_{\mu}(k)$ departs with the decomposition into its real and imaginary parts known from [14, p. 1877] and is repeated here for convenience. First observe
@@
by expanding (4.24) with $H_{\mu}^{(2)}(k)=J_{\mu}(k)-\operatorname{i}Y_{\mu}(k)$. The Wronskian identity $J_{\mu}(k)Y_{\mu}^{\prime}(k)-Y_{\mu}(k)J_{\mu}^{\prime}(k)=2/(\pi k)$ from [4, §10.5.2] is a consequence of the recurrence relations for Bessel functions. This reveals the characterisation [14, Eqn. (3.10)] of $w_{\mu}$, namely
@@
Since $kM_{\mu}^{2}(k)$ is non-increasing for $\mu\geq 1/2$ and converges to $2/\pi$ as $k\rightarrow\infty$ [19, §13.74], one has $2\leq\pi kM_{\mu}^{2}(k)<\infty$ and
@@
The situation is different for $\mu=0$ where $M_{0}^{2}(k)$ admits no lower bound for $k\rightarrow 0$ that is linear in $1/k$. However, the asymptotic expansion [19, §13.75] of $M_{0}^{2}(k)$ reveals the convergence $kM_{0}^{2}(k)\rightarrow 2/\pi$ as $k\rightarrow\infty$ with the lower bound (see [4, §10.18.iii])
@@ 4.3 Spectral representation of the Laplace DtN
4.4 Proofs of Theorems 3.4 and 3.5
The sharp bounds in Theorem 4.1 and the explicit representation of the DtN operator (4.13) enable the proofs of Theorems 3.4 and 3.5 stated in Subsection 3.2.
Proof of Theorem 3.4.
@@
$\mathopen{}\mathclose{{}\left\langle\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)g,\overline{h}}\right\rangle_{S_{1}}=\sum_{m\in\mathbb{N}_{0}}\sum_{j\in J_{m}}z_{m,\nu}(s)\,\widehat{g}_{m,j}\,\overline{\widehat{h}_{m,j}}=\mathopen{}\mathclose{{}\left\langle g,\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)\overline{h}}\right\rangle.$ (4.41)
Since $z_{m,\nu}(s)=\overline{z_{m,\nu}(\overline{s})}$ by Theorem 4.1.i, this proves Theorem 3.4.i. Recall from the definition of the Fourier coefficients (4.11) and the equivalence of norms [11, Rem. 7.6] in $H^{1/2}(S_{1})$ that there exists a constant $C_{1}(n)>0$ that exclusively depends on $n$ such that
@@
Proof of Theorem 3.5.
Recall the Hilbert space $H^{1}(\Omega,S_{R})$ from (2.5) (for $\Gamma=S_{R}$).
Assume for a contradiction to (3.6) that there exists two sequences $(v_{k})_{k\in\mathbb{N}}\subset H^{1}(\Omega,S_{R})$ and $(s_{k})_{k\in\mathbb{N}}\subset\mathbb{C}_{\geq 0}$ with $\partial_{r}v_{k}=\operatorname{DtN}(s_{k})v_{k}$ on $S_{R}$ such that $\mathopen{}\mathclose{{}\left\|v_{k}}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}=1$ for all $k\in\mathbb{N}$ and $\|\nabla v_{k}\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}+\mathopen{}\mathclose{{}\left\|\partial_{r}v_{k}}\right\|_{H^{-1/2}(S_{R})}\rightarrow 0$ as $k\rightarrow\infty$. Since the sequence $(v_{k})_{k\in\mathbb{N}}$ is bounded in $H^{1}(\Omega,S_{R})$, it admits a weakly convergent subsequence with weak limit $u\in H^{1}(\Omega,S_{R})$. A standard compactness argument with the compact embedding $H^{1}\mathopen{}\mathclose{{}\left(\Omega}\right)\hookrightarrow L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)$ and the weak lower semicontinuity of norms reveal $\mathopen{}\mathclose{{}\left\|u}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}=1$ and $\mathopen{}\mathclose{{}\left\|\nabla u}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}=0=\mathopen{}\mathclose{{}\left\|\partial_{r}u}\right\|_{H^{-1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right)}$. Hence $u\in H^{1}(\Omega,S_{R})$ is constant on the (in particular connected) domain $\Omega$.
The traces of $v_{k}$ and the constant $u$ on $S_{R}$ have the Fourier expansions
$v_{k}|_{S_{R}}=\sum_{m\in\mathbb{N}_{0}}\sum_{j\in J_{m}}\widehat{v}_{m,j}^{(k)}Y_{m,j}\qquad\text{and}\qquad u|_{S_{R}}=\widehat{u}_{0,0}Y_{0,0}\quad\text{for }\widehat{u}_{0,0}=\frac{u}{Y_{0,0}}=\frac{u}{|S_{R}|^{1/2}}$
with $\widehat{v}_{\bullet,\bullet}^{(k)}\in\mathbb{C}$ for all $k\in\mathbb{N}$ in terms of the $L^{2}(S_{R})$-orthonormal Laplace-Beltrami eigenfunctions $Y_{\bullet,\bullet}$ from (4.1). The DtN operator representation (4.13) and $\partial_{r}v_{k}=\operatorname{DtN}(s_{k})v_{k}$ result in
$\frac{z_{0,\nu}(s_{k}R)}{R}\widehat{v}_{0,0}^{(k)}\,\overline{\widehat{u}_{0,0}}=\mathopen{}\mathclose{{}\left\langle\operatorname{DtN}(s_{k})v_{k},\overline{u}}\right\rangle_{S_{R}}=\mathopen{}\mathclose{{}\left\langle\partial_{r}v_{k},\overline{u}}\right\rangle_{S_{R}}\to 0$
with the weak convergence $\partial_{r}v_{k}\rightharpoonup\partial_{r}u=0$ in $H^{-1/2}(S_{R})$ as $k\to\infty$ in the last step. This and $1\leq\nu+1/2\leq|z_{0,\nu}(s_{k}R)|$ for $n\geq 3$ from Theorem 4.1.ii establish
$|\widehat{v}_{0,0}^{(k)}||\widehat{u}_{0,0}|\leq|z_{0,\nu}(s_{k}R)||\widehat{v}_{0,0}^{(k)}||\widehat{u}_{0,0}|=R\,\mathopen{}\mathclose{{}\left|\mathopen{}\mathclose{{}\left\langle\operatorname{DtN}(s_{k})v_{k},\overline{u}}\right\rangle_{S_{R}}}\right|\to 0\qquad\text{as }k\to\infty.$
Thus either $\widehat{u}_{0,0}=0$ or $\widehat{v}_{0,0}^{(k)}\to 0$ as $k\to\infty$ hold. The weak convergence $v_{k}|_{S_{R}}\rightharpoonup u|_{S_{R}}$ in $H^{1/2}(S_{R})$ (by the continuity of the trace map) implies the convergence $\widehat{v}_{0,0}^{(k)}\to\widehat{u}_{0,0}$ in $\mathbb{C}$ as $k\to\infty$ and reveals $\widehat{u}_{0,0}=0$ in the second case. Hence $u=0$ follows in any case. This contradicts $\|u\|_{L^{2}(\Omega)}=1$ and concludes the proof of (3.6). ∎
@@
The proof of Proposition 3.2 follows mainly from the theory in [13], where integral equation techniques are used. The key role is played by Green’s representation formula and we sketch the arguments in the following. The starting point is the exterior problem [13, Eqn. (7.31)] ( for $\mathcal{P}=-\Delta+s^{2}$, $f=0$, and $\Gamma=S_{R}$ therein) written as:
@@
Here $\mathcal{M}u=0$ is a condition for the behaviour at infinity which is defined in [13, (7.29)]. Theorems 8.9 and 9.6 in [13] translate this condition for our setting to
@@
If the exterior problem (3.1) with condition $\operatorname*{IC}\nolimits_{s}^{\operatorname*{Green}}$ has a solution $u$, then this solution satisfies Green’s representation theorem by [13, Thm. 7.15], namely
@@
where $\operatorname*{SL}\mathopen{}\mathclose{{}\left(s}\right)$ and $\operatorname*{DL}\mathopen{}\mathclose{{}\left(s}\right)$ are the single and double layer potential operators for the operator $\mathopen{}\mathclose{{}\left(-\Delta+s^{2}}\right)$ (see, e.g., [13, Eqn. (7.1)] for details) and $\psi$ is a solution of
@@
with the abbreviations $S:=\frac{1}{2}\mathopen{}\mathclose{{}\left(\gamma^{+}\operatorname*{SL}\mathopen{}\mathclose{{}\left(s}\right)+\gamma^{-}\operatorname*{SL}\mathopen{}\mathclose{{}\left(s}\right)}\right)$ and $T:=\frac{1}{2}\mathopen{}\mathclose{{}\left(\gamma^{+}\operatorname*{DL}\mathopen{}\mathclose{{}\left(s}\right)+\gamma^{-}\operatorname*{DL}\mathopen{}\mathclose{{}\left(s}\right)}\right)$. Note that for any $g\in H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right)$ and $h\in H^{-1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right)$, $\operatorname*{DL}\mathopen{}\mathclose{{}\left(s}\right)g$ and $\operatorname*{SL}\mathopen{}\mathclose{{}\left(s}\right)h$ define functions in $\mathbb{R}^{n}\backslash S_{R}$ and $\gamma^{+}$ denotes the trace from the outer domain $B_{R}^{+}$, while $\gamma^{-}$ is the trace from the inner domain $B_{R}$. It is proved in [13, Theorem 6.11] that the mappings
@@
are continuous linear operators. In this way, well-posedness of (3.1) with condition $\operatorname*{IC}\nolimits_{s}^{\operatorname*{Green}}$ follows if $S$ in (A.1) has a bounded inverse. Theorem 7.6 in [13] implies that $S$ is a Fredholm operator of index 00 so that uniqueness implies bounded invertibility. Uniqueness follows for $n\geq 2$ and $s\in\mathbb{C}^{\bullet}_{\geq 0}$ from [13, Theorem 9.11], for $n\geq 3$ and $s=0$ from [13, Corollary 8.11] and for $n=2$, $s=0$, and $R\neq 1$ from [13, Theorem 8.16.(ii)]. Hence (3.1) with condition $\operatorname*{IC}\nolimits_{s}^{\operatorname*{Green}}$ is well-posed in these three cases (with continuous solution operator). This finishes the proof of Proposition 3.2.1a and 3.2.2b, whereas 3.2.1b directly follows from the implication (4.10).
It remains to prove Proposition 3.2.2a. Uniqueness of solutions is guaranteed by [13, Thm 8.10]. Existence and well-posedness follows from the expansion (4.38) of the solution